# Next-Generation Reservoir Computing for Dynamical Inference

### v1

@@ IV.1.1 Lorenz system
Figure 4 shows the true Lorenz trajectory (test data) alongside the trajectory generated by the reservoir model using iterative one-step predictions. As expected for a chaotic system, the two trajectories diverge after some time, where small differences amplify exponentially in time. The divergence occurs on a timescale comparable to the Lyapunov time due to the finite accuracy of our one-step predictor.
Figure 4: Short-term trajectories of the Lorenz system. The true trajectory is shown in blue, and the reservoir model prediction in orange. The model was trained using only the scalar variable $x(t)$.
@@
Figure 5: Delay-embedded Lorenz attractor reconstruction. The true attractor is shown on the left, and the reconstructed attractor on the right. The model was trained using only the scalar variable $x(t)$.
@@ IV.1.2 Rössler system
Figure 6: Short-term trajectories of the Rössler system. The true trajectory is shown in blue, and the reservoir model prediction in orange. The model was trained using only the scalar variable $x(t)$.
@@
Figure 7: Delay-embedded Rössler attractor reconstruction. The true attractor is shown on the left, and the reconstructed attractor on the right. The model was trained using only the scalar variable $x(t)$.
@@ IV.2 Reconstruction of bifurcation diagram
We consider the Rössler system (11) and add a variable controll parameter $\eta$ to the right-hand-side of the $y$ equation. The reservoir then ovserves the resulting system states and learns to repdouce the system’s behaviour for different $\eta$. The variation of $\eta$ allows us to probe the bifurcation structure of the system. In the following, we first use several data sections corresponding to different constant $\eta$ values to reconstruct the bifurcation diagram, and secondly, consider a time-dependent, stochastic $\eta(t)$ to mimic less controlled real world situations.
@@
The reconstruction is accurate near the trained parameter values, and is well interpolated between them. As $\eta$ moves farther away from $\eta$ values used for training, the reservoir-generated attractors increasingly deviate from the true system attractor. This is expected, since the model extrapolates beyond the regimes it has seen during training. In Fig.8, this is most prominent near $\eta=-1$, where the reservoir predicts qualitatively different behavior [36].
@@
Next, we consider a more challenging and practically relevant scenario. In many real-world applications, it is not possible to collect data under strictly controlled conditions where the control parameter (here, $\eta$) remains constant. Instead, the parameter often varies in time and can be measured only as a noisy signal.
To mimic this situation, we drive the Rössler system  (11) with a stochastic parameter $\eta(t)$ evolving according to an Ornstein-Uhlenbeck (OU) process [37]:
@@
$\dot{\eta}$ $=\frac{1}{\tau}\left(-\eta+\rho\xi\right)$
where $\xi(t)$ is Gaussian white noise, and the parameters $\tau$ and $\rho$ control the correlation time and amplitude of the OU process, respectively. The resulting system has a time-varying, stochastic control parameter, making the learning task harder. We use the trajectories generated by this noise-driven system as training data for a single reservoir model.
In this case as well, the reservoir accurately reproduces the bifurcation structure near trained values of $\eta$, demonstrating that even from trajectories with rapidly time-varying control parameters, it can learn the correct asymptotic dynamics corresponding to fixed parameter values. Deviations arise primarily when extrapolating beyond the training range, see Figure 9.
Figure 9: Top panel: Bifurcation diagram of the true Rössler system (11) (blue) and the reconstruction from the reservoir (orange). The training data were obtained from a Rössler system perturbed by an Ornstein-Uhlenbeck input (Eq. (12)). The green curve shows the empirical density of the noise variable $\eta$. The bifurcation points are computed by plotting the local maxima of $x(t)$. Training used a long trajectory ($T=10^{5}$), all three variables ($x,y,z$), reservoir size $M=6000$, delay dimension $H=25$, and 1% measurement noise. Bottom panel: The training signal, Rössler variables $(x,y,z)$ are shown in blue, cyan, and purple, and the OU input $\eta(t)$ in green (vertically shifted for clarity).
@@ V Discussion
In our work, we built on the next-generation reservoir computing framework (NGRC) and proposed a projection method based on pseudorandom nonlinear features. An advantage of this approach is that the number of projected features $M$ can be set independently of the input dimension, whereas in polynomial-based projections the feature space typically depends on the input size. We demonstrated the utility of this scalable scheme across several canonical examples, including attractor reconstruction, bifurcation diagram estimation, and asymptotic phase inference. The models showed robust performance even when trained on limited or non-stationary data, and they could generalize to parameters and dynamics not seen during training. These features suggest that next-generation reservoir computing may be a practical tool for applications ranging from digital twins and surrogate modeling to problems in climate science, finance, and biology.
@@
A known challenge in NGRC, highlighted in Ref. [32], is the risk of transversal instability, where small perturbations orthogonal to the attractor direction can grow, degrading long-term predictions. Interestingly, we did not observe such instabilities in any of the examples tested. The models remained stable over long trajectories, including during bifurcation diagram estimation via quasi-continuation. Although we did not conduct a systematic study of this effect, we speculate that adding small amounts of measurement noise during training acts as a form of implicit regularization and may contribute to improved transversal stability.
@@ Acknowledgements.
We thankfully acknowledge useful discussions with Andrew Flynn and Herbert Jaeger.

### v2

@@ IV.1.1 Lorenz system
Figure 6 shows the true Lorenz trajectory (test data) alongside the trajectory generated by the reservoir model using iterative one-step predictions. As expected for a chaotic system, the two trajectories diverge from each other after some time, where small differences amplify exponentially in time. This divergence occurs on a timescale comparable to the Lyapunov time due to the finite accuracy of our one-step predictor.
Figure 6: Short-term trajectories of the Lorenz system. The true trajectory is shown in blue, and the reservoir model prediction in orange. The model was trained using only the scalar variable $x(t)$. Time axis is units of time-sampling step $\tau=0.04$.
@@
Figure 7: Delay-embedded Lorenz attractor reconstruction. The true attractor is shown on the left, and the reconstructed attractor on the right. The model was trained using only the scalar variable $x(t)$. Time delay for plot is $3\tau=0.12$.
@@ IV.1.2 Rössler system
Figure 8: Short-term trajectories of the Rössler system. The true trajectory is shown in blue, and the reservoir model prediction in orange. The model was trained using only the scalar variable $x(t)$. Time axis is units of time-sampling step $\tau=0.15$.
@@
Figure 9: Delay-embedded Rössler attractor reconstruction. The true attractor is shown on the left, and the reconstructed attractor on the right. The model was trained using only the scalar variable $x(t)$. Time delay for plot is $8\tau=1.2$.
@@ IV.2 Reconstruction of bifurcation diagram
We consider the Rössler system (12) and add a variable control parameter $\eta$ to the right-hand-side of the $y$ equation. The reservoir then observes the resulting system states and learns to reproduce the system’s behaviour for different $\eta$. The variation of $\eta$ allows us to probe the bifurcation structure of the system. In the following, we first use several data sections corresponding to different constant $\eta$ values to reconstruct the bifurcation diagram, and secondly, consider a time-dependent, stochastic $\eta(t)$ to mimic less controlled real world situations.
@@
The reconstruction is accurate near the trained parameter values, and is well interpolated between them. As $\eta$ moves farther away from $\eta$ values used for training, the reservoir-generated attractors increasingly deviate from the true system attractor. This is expected, since the model extrapolates beyond the regimes it has seen during training. Such deviations are a common limitation of data-driven models that rely on local interpolation rather than a global parameterization of the underlying dynamics. In particular, NGRC has no built-in mechanism to guarantee correct dynamics outside the region explored during training, and uncertainties can grow rapidly once the model leaves this domain. In Fig. 10, this is most prominent near $\eta=-1$, where the reservoir predicts qualitatively different behavior [44].
@@
Next, we consider a more challenging and practically relevant scenario. In many real-world applications, it is not possible to collect data under strictly controlled conditions where the control parameter, $\eta$, remains constant. Instead, the parameter often varies in time and can be measured only as a noisy signal.
To mimic this situation, we drive the Rössler system  (12) with a stochastic parameter $\eta(t)$ evolving according to an Ornstein-Uhlenbeck (OU) process [45]. The OU process is a standard way to model noise with a finite correlation time. It describes a fluctuating variable that is randomly driven but does not change abruptly, producing smoothly varying noise rather than instantaneous (white-noise) jumps.
@@
$\dot{\eta}$ $=\frac{1}{\tau_{c}}\left(-\eta+\rho\xi\right)$
where $\xi(t)$ is Gaussian white noise, and the parameters $\tau_{c}$ and $\rho$ control the correlation time and amplitude of the OU process, respectively. The resulting system has a time-varying, stochastic control parameter, making the learning task harder. We use the trajectories generated by this noise-driven system as training data for a single reservoir model.
In this case as well, the reservoir accurately reproduces the bifurcation structure near trained values of $\eta$, demonstrating that even from trajectories with rapidly time-varying control parameters, it can learn the correct asymptotic dynamics corresponding to fixed parameter values. Deviations arise primarily when extrapolating beyond the training range, see Fig. 11. This again reflects the fact that the model is asked to infer behavior for parameter values that are only sparsely or indirectly represented in the training data.
Figure 11: Top panel: Bifurcation diagram of the true Rössler system (12) (blue) and the reconstruction from the reservoir (orange). The training data were obtained from a Rössler system perturbed by an Ornstein-Uhlenbeck input (Eq. (13)). The green curve shows the empirical probability density of the noise variable $\eta$. The bifurcation points are computed by plotting the local maxima of $x(t)$. Training used a long trajectory ($T=10^{5}$), all three variables ($x,y,z$), reservoir size $M=6000$, delay dimension $H=25$, and 1% measurement noise. Bottom panel: The training signal, Rössler variables $(x,y,z)$ are shown in blue, cyan, and purple, and the OU input $\eta(t)$ in green (vertically shifted for clarity).
@@ V Discussion
In our work, we built on the next-generation reservoir computing framework (NGRC) and proposed a projection method based on pseudorandom nonlinear features. An advantage of this approach is that the number of projected features $M$ can be set independently of the input dimension, whereas in polynomial-based projections the feature space typically depends on the input size $N\cdot H$. We demonstrated the utility of this scalable scheme across several canonical examples, including attractor reconstruction, bifurcation diagram estimation, and asymptotic phase inference. The models showed robust performance even when trained on limited or non-stationary data, and they could generalize to parameters and dynamics not seen during training. These features suggest that next-generation reservoir computing may be a practical tool for applications ranging from digital twins and surrogate modeling to problems in climate science, finance, and biology. Reservoir computing has also been applied to real-world forecasting problems, such as anticipating food price fluctuations in agricultural markets [46].
@@
A known challenge in NGRC, highlighted in Ref. [34], is transversal instability in long-term autonomous prediction, whereby small perturbations orthogonal to the attractor can grow and lead to rapid divergence. We examined this issue directly and found that, although Tikhonov regularization helps mitigate overfitting and numerical ill-conditioning in one-step regression — consistent with conditioning effects reported in Ref. [38] — it does not prevent long-term instabilities: models trained without noise generally diverge quickly when iterated autonomously, even when regularized. By contrast, adding a small amount of measurement noise to the training inputs reliably suppresses such failures (similar stabilizing effects were reported in classical RC [41]). Post-processing approaches for enhancing long-horizon robustness have also been considered [40]. Across all systems tested, we did not observe any transverse instability once 1% measurement noise was included, even for extremely long autonomous rollouts used in our quasi-continuation bifurcation reconstructions. A plausible explanation is that noise spreads the training samples slightly off the attractor surface, so the regression must learn a mapping that remains locally consistent not only along the attractor but also in nearby transverse directions (see also the structural stability considerations discussed in Ref. [39]). Deterministic trajectories sample only a thin manifold and leave transverse directions weakly constrained; noise broadens this sampling, which empirically yields models that do not amplify small off-manifold deviations. We therefore use 1% input noise as our default configuration in all experiments.
@@ Acknowledgements.
We thankfully acknowledge useful discussions with Andrew Flynn and Herbert Jaeger. We acknowledge financial support from the Crafoord Foundation (project no. 20240689).